please help

Schopp Company makes swimsuits and sells these suits directly to retailers. Although Schopp has a variety of suits, it does not make the All-Body suit used by highlyskilled swimmers. The market research department believes that a strong market exists for this type of suit. The department indicates that the All-Body suit would sellfor approximately $114. Given its experience, Schopp believes the All-Body suit would have the following manufacturing costs.
Direct materials $29
Direct labor 28
Manufacturing overhead 45
Total costs $102

Assume that Schopp uses cost-plus pricing, setting the selling price 24% above its costs. What would be the price charged for the All-Body swimsuit?


assume that Schopp uses target costing. What is the price that Schopp would charge the retailer for the All-Body swimsuit?

What is the highest acceptable manufacturing cost Schopp would be willing to incur to produce the All-Body swimsuit, if it desired a profit of $26 per unit?
Answer #1
Concepts and reason

Cost Accounting: Cost accounting refers to the process of recording, classifying, summarizing, allocating and analyzing the different alternative measures to control costs. Its purpose is to guide the management on how to control costs effectively and maximize productivity. It provides detailed information of the costs that needs to be focused and find out the alternative course of actions to control the costs in the future.

Types of Cost Accounting:

• It is a method of alternating the estimated cost with an actual cost and showing the periodic variances of the estimated cost with its actual costs.

Standard cost accounting:

• Activity-based costing: It is a method of identifying the activities performed by a company and allocates the indirect costs to the products and services based on the actual consumption by each product or service.

• Resource consumption accounting: It is a management theory that explains the comprehensive management accounting approach which helps the management to make decisions for the company’s overall growth.

• Throughput accounting: It is an accounting approach which helps the management to make decisions for the company’s profitability growth.

• Life Cycle Costing: It is a measure used to analyze the pending and future costs and find out alternative measure that would have a positive impact on such costs. It compares the initial investment possibilities and evaluates the least cost courses of actions for a period of twenty years.

• Environmental Accounting: It is an accounting that identifies the use of natural resources, analyzes the costs associated with it and its impact on the environment and provides the detailed cost information with alternative course of actions.

• Target Costing: It is an accounting method to evaluate the cost of product’s life cycle which would be effective, functional, qualitative and able to generate expected profit.

Fundamentals

Cost plus pricing: Cost plus pricing is a method in which selling price is calculated by adding direct materials, direct labor and manufacturing overhead for a product and then adding specific markup cost to a product unit cost. It is otherwise called as value based pricing.

Formula for price charged for All-Body swimsuit is shown below:

PricechargedforAllBodyswimsuit=Totalcostsperunit+Profitperunit{\rm{Price charged for All - Body swimsuit}}\,{\rm{ = }}\,{\rm{Total costs}}\,{\rm{per unit + Profit}}\,{\rm{per unit}}

Total Costs: Total cost is the sum of fixed, variable and mixed expenses incurred to produce a product. While the fixed costs remain unchanged (neither increase nor decrease) with quantity of goods produced, variable costs increase with the increase in quantity of production and reduce with the decrease in production.

Selling price: The term ‘selling price’ is defined as the price at which a good or service is sold by the seller to the buyer. It is generally expressed as currency units. In other words, it is a market value or agreed exchange value that enables a buyer to purchase goods or services. It is also known as list price, quoted price, market price, or sale price.

Formula for highest acceptable manufacturing cost per unit as shown below:

HighestacceptablemanufacturingcostperunitofSCompany]=SellingpriceperunitDesiredprofitperunit\left. \begin{array}{l}\\{\rm{Highest acceptable manufacturing }}\\\\{\rm{cost per unit of S Company}}\\\end{array} \right]\,{\rm{ = }}\,{\rm{Selling price per unit }} - \,{\rm{Desired profit}}\,{\rm{per unit}}

Calculate price charged for the All-Body swimsuit of S Company.

PricechargedforAllBodyswimsuit=Totalcostsperunit+Profitperunit=$102+$24.48=$126.48Perunit\begin{array}{c}\\{\rm{Price charged for All - Body swimsuit}}\,{\rm{ = }}\,{\rm{Total costs}}\,{\rm{per unit + Profit}}\,{\rm{per unit}}\\\\{\rm{ = }}\,{\rm{\$ 102}}\,{\rm{ + }}\,{\rm{\$ 24}}{\rm{.48}}\\\\{\rm{ = }}\,{\rm{\$ 126}}{\rm{.48}}\,{\rm{Per unit}}\\\end{array}

Therefore, price charged for All-Body swimsuit of S Company is $126.48 per unit.

Working note:

Calculation of Profit for All-Body swimsuit of S Company as shown below:

ProfitforAllBodyswimsuitofSCompany=TotalCostsperunit×PercentofTotalCosts=$102×24%=$24.48Perunit\begin{array}{c}\\{\rm{Profit for All - Body swimsuit of S Company}}\, = \,{\rm{Total Costs}}\,{\rm{per unit}}\, \times \,{\rm{Percent of Total Costs}}\\\\{\rm{ = }}\,{\rm{\$ 102}}\, \times \,{\rm{24\% }}\\\\{\rm{ = }}\,{\rm{\$ 24}}{\rm{.48}}\,{\rm{Per unit}}\\\end{array}

Hence, profit for All-Body swimsuit of S Company is $24.48 per unit.

It is given that All-Body swimsuit would be sold for $114. Hence, the price that S Company would charge the retailer for all Body swimsuits is $114 per unit.

Calculate highest acceptable manufacturing cost per unit of S Company.

HighestacceptablemanufacturingcostperunitofSCompany]=SellingpriceperunitDesiredprofitperunit=$114$26=$88Perunit\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Highest acceptable manufacturing }}\\\\{\rm{cost per unit of S Company}}\\\end{array} \right]\,{\rm{ = }}\,{\rm{Selling price per unit }} - \,{\rm{Desired profit}}\,{\rm{per unit}}\\\\{\rm{ = }}\,{\rm{\$ 114}}\, - \,\$ 26\\\\ = \,\$ 88\,{\rm{Per unit}}\\\end{array}

Therefore, highest acceptable manufacturing cost per unit of S Company is $88 per unit.

Ans:

Price charged for All-Body swimsuit of S Company is $126.48 per unit.

The price that S Company would charge the retailer for all Body swimsuits is $114 per unit.

Highest acceptable manufacturing cost per unit of S Company is $88 per unit.